a car moving at 10m/s, accelerated uniformely over a distance of 25m, at the end of which the velocity of the car became 15m/s. determine the acceleration of the car and the time taken to cover the 25m
1
Expert's answer
2020-04-21T19:12:36-0400

Solution.

"\upsilon_0=10 m/s;"

"S=25m;"

"\upsilon=15 m/s;"

"a - ?; t - ?"

"S = \dfrac{\upsilon^2 -\upsilon_0^2 }{2a}; \implies a=\dfrac{\upsilon^2 -\upsilon_0^2 }{2S};"

"a=\dfrac{(15 m/s)^2 -(10 m/s)^2 }{2\sdot25 m}=2.5 m/s^2;"

"S = \dfrac{\upsilon +\upsilon_0 }{2}\sdot t \implies t = \dfrac{2\sdot S }{\upsilon +\upsilon_0};"

"t = \dfrac{2\sdot25 m }{ 15 m/s+10 m/s}=2s;"

Answer:"a=2.5 m\/s^2;\nt = 2s."